# Automated Trustworthiness Oracle Generation for Machine Learning Text Classifiers

### v1

@@ 6.5. Results
We use model confidence as a metric to assess the uncertainty of ML predictions. Figure 6 addresses RQ1.1 by illustrating the distribution of model confidence for two trustworthiness labels on the benchmark shown in Table 2. The majority of high confidence predictions are trustworthy. Highly confident (0.9–1.0) predictions that are trustworthy account for 76%–94% of predictions with highly strong confidence. Similarly, trustworthy predictions with high confidence (0.8–0.9) represent 49%–91% of high confidence predictions. While predictions with strong confidence are likely to be trustworthy, several confident predictions are still untrustworthy. Confident (0.8–1.0) but untrustworthy predictions make up 6%–29% of all confident predictions. In contrast, predictions with low confidence (0.8) can also be trustworthy, with 51%–91% of low confidence predictions being trustworthy.
@@
We now address RQ1.2, finding the optimal value of conf for Naive, which relies solely on model confidence. Figure 9 shows receiver-operating characteristic (ROC) curves of Naive across trustworthiness datasets. All ROC curves are near the random curve, indicating model confidence is limited to distinguish between trustworthy and untrustworthy predictions. Figure 6 further shows that increasing conf on the y-axis decreases the false positive rate, where untrustworthy predictions are labeled as trustworthy, but also increases the false negative rate, where trustworthy predictions are labeled as untrustworthy. In other words, increasing conf enhances specificity but reduces sensitivity. Hence, we choose to maximise G-mean as the criteria to find conf. Figure 9 illustrates Naive’s effectiveness in G-mean with different values of conf. As the value of 0.9 achieves the best G-mean for Naive, we set conf to 0.9 for the remaining experiments.
![[Uncaptioned image]](x15.png)
@@
We also address RQ2.1 by measuring the effectiveness of TOKI-lime in G-mean with various values of dist. The result shown in Figure 9 reveals that the optimal value of dist ranges from 0.2 to 0.4, depending on the models under test. Across all four trustworthiness datasets, when dist exceeds 0.6, G-mean drops significantly. In this case, TOKI considers all words as keywords and biases for labeling predictions as trustworthy. In the subsequent experiments, we set dist to 0.3, as it balances the effectiveness across the benchmark.
@@
RQ2.2: Comparing TOKI and Naive. TOKI consistently outperforms Naive across all datasets. Naive shows good precision but relatively low specificity and G-mean, especially on Dong’s and our trustworthiness datasets. This indicates that Naive is biased toward labeling predictions as trustworthy, resulting in misclassifying untrustworthy predictions as trustworthy.
@@
RQ2.4: The impact of explanation methods. Overall, TOKI demonstrates the best effectiveness with the state-of-the-art LIME. TOKI-lime achieves higher accuracy and G-mean compared to TOKI-omis and TOKI-grad, while other metrics show no significant differences, except for the CAMS dataset. Explanation methods also impact TOKI’s efficiency. TOKI-lime is significantly slower due to LIME’s sampling of 5,000 neighbors, while omission and gradient are more lightweight. The model complexity also affects TOKI’s efficiency, as the fine-tuned models consume more time than the two simpler MLPs in Dong’s dataset.
@@
Table 4 compares the effectiveness of TOKI-guided attack method and A2T (Yoo and Qi, 2021). Overall, TOKI-guided attack method outperforms A2T, achieving a higher ASR and lower NP across all models. Adversarial examples generated by TOKI also have higher Bert and USE scores than those by A2T, except for the USE score on the ag_news dataset, demonstrating TOKI’s ability to preserve the semantic meaning of the original texts. Table 5 displays several adversarial examples generated by both methods. It is observed that TOKI can effectively find synonyms to successfully attack with fewer perturbed words, while A2T requires more perturbations and sometimes fails.
[Answer to RQ3] The adversarial attack method guided by TOKI outperforms the state-of-the-art A2T. Our method achieves a higher success rate with fewer perturbations than A2T while preserving the semantic meaning of the original texts.
@@ 7.1. Implications
Experimental results show that highly confident predictions are more likely to be trustworthy. However, high confidence does not guarantee trustworthiness, and low confidence predictions can still be trustworthy. Relying on prediction uncertainty is ineffective for trustworthiness assessment, overlooking both trustworthy low and untrustworthy high confidence predictions. Prior work has a complex debate over the impact of prediction uncertainty, mainly on human trust in ML models. Several studies (Nguyen et al., 2015; Bussone et al., 2015) suggest that prediction uncertainty has a limited impact on human trust (Rechkemmer and Yin, 2022; Ovadia et al., 2019). Other studies (Zhang et al., 2020a) show that prediction certainty can improve human trust in ML models and increase the willingness to rely on high confidence predictions. However, we argue that whether the certainty of a model’s predictions increases human trust or not, it does not translate into improving the model’s trustworthiness. Model certainty can lead to overreliance on ML decisions without reflecting the true trustworthiness of the model. The model may make decisions based on spurious correlations, being confident in them rather than valid correlations in the training data. This fits the distinction between trust and trustworthiness defined by Kästner et al. (Kästner et al., 2021), where people can still trust an untrustworthy model.
In contrast, TOKI shows its superior effectiveness compared to the baselines. TOKI identifies keywords for each class and uses them to measure the semantic relatedness between words and the class. This approach better captures the characteristics of classes, outperforming its variant that does not use keywords to measure relatedness, which is biased for untrustworthy predictions.
@@ 7.2. Threats to Validity
To reduce threats to external validity, we evaluate TOKI on approximately 6,000 predictions across 10 models and datasets in various domains, including topic classification, sentiment analysis, clinical mental text classification, and hate speech detection. We also employ TOKI with various explanation methods to assess its performance.
Threats to construct validity can arise from the trustworthiness benchmark. We derive trustworthiness datasets from existing studies, which do not directly label predictions as trustworthy or not, by measuring the explanations’ plausibility. To mitigate the threats, we collect an additional, more straightforward trustworthiness dataset. The bias in human annotations is addressed by combining all human-annotated trustworthiness labels for each prediction using plurality voting (Leon et al., 2017).
@@ 8.2. Making Use of Explanations
Explanations have served as a valuable tool for testing and improving ML models (Zhao et al., 2024). Various studies have used explanations to understand and debug models. Ribeiro et al. (Ribeiro et al., 2016) demonstrated the usefulness of local explanations in various tasks, such as model comparison and trust assessment. Lapuschkin et al. (Lapuschkin et al., 2019) introduced a semi-automated approach that characterises and validates classification strategies. Thomas et al. (Thomas et al., 2019) used explanations to uncover input-prediction patterns. Several studies (Yoo and Qi, 2021; Zhang et al., 2020b; Chai et al., 2023) have applied explanations in adversarial attacks, primarily to determine word substitution order based on word importance. To the best of our knowledge, this paper introduces the first approach that leverages explanations to make perturbations, specifically, by weakening valid correlations and strengthening spurious ones. Explanations have also been integrated into the learning process to improve model performance and reliability. Andrew et al. (Ross et al., 2017) proposed an approach to discover multiple models for the same task with different classification strategies, allowing domain experts to choose the best one. Chen et al. (Chen and Ji, 2022) used explanations to improve the adversarial robustness of language models. Other studies (Schramowski et al., 2020; Ghai et al., 2021) leveraged explanations to make models right for the right scientific reasons. Similarly, Du et al. (Du et al., 2021) developed a framework to mitigate shortcuts focusing on stop words, punctuation, and numbers. Recent studies (Linhardt et al., 2024) have also applied explanation regularisation to mitigate the impact of spurious correlations.
@@ 9. Conclusion
This paper investigates the trustworthiness oracle problem of text classifiers. Statistical evaluation reveals that while highly confident predictions are more likely to be trustworthy, some still lack trustworthiness due to reliance on spurious correlations. We propose TOKI, an automated trustworthiness oracle generation method. Experiments are conducted to compare the effectiveness and efficiency of TOKI, TOKI’s ablation variant and the naive approach against the trustworthiness benchmark on ten models and datasets. Results show that TOKI outperforms other approaches. We also introduce a TOKI-guided adversarial attack method, which proves to be more effective than the state-of-the-art A2T. In addition, several works can be explored in the future, including adapting TOKI for other data types and conducting a larger human-based evaluation to gain more insights into the trustworthiness problem.
Data Availability
The trustworthiness benchmark and the replicate package are available on Zenodo (Anonymous, 2024).

### v3

@@ 6.5. Results
We use model confidence as a metric to assess the uncertainty of ML predictions. Figure 6 addresses RQ1.1 by illustrating the distribution of model confidence for two trustworthiness labels on the benchmark shown in Table 2. The majority of high confidence predictions are trustworthy. Highly confident (0.9–1.0) predictions that are trustworthy account for 76%–99% of predictions with highly strong confidence. Similarly, trustworthy predictions with high confidence (0.8–0.9) represent 49%–91% of high confidence predictions. While predictions with strong confidence are likely to be trustworthy, several confident predictions are still untrustworthy. Confident (0.8–1.0) but untrustworthy predictions make up 2%–29% of all confident predictions. In contrast, predictions with low confidence (0.8) can also be trustworthy, with 51%–92% of low confidence predictions being trustworthy.
@@
![[Uncaptioned image]](x17.png)
@@
We now address RQ1.2, finding the optimal value of conf for Naive, which relies solely on model confidence. Figure 9 shows receiver-operating characteristic (ROC) curves of Naive across trustworthiness datasets. Only the ROC curve for Issues is slightly closer to the top-left corner of the plot, while the others are near the random curve. Notably, the ROC curve for the new trustworthiness dataset falls below the random curve. This indicates that model confidence is limited in distinguishing between trustworthy and untrustworthy predictions. Figure 6 further shows that increasing conf on the y-axis decreases the false positive rate, where untrustworthy predictions are labeled as trustworthy, but also increases the false negative rate, where trustworthy predictions are labeled as untrustworthy. In other words, increasing conf enhances specificity but reduces sensitivity. Hence, we choose to maximise G-mean as the criterion to find conf. Figure 9 illustrates Naive’s effectiveness in G-mean with different values of conf. As the value of 0.9 achieves the best G-mean for Naive, we set conf to 0.9 for the remaining experiments.
@@
Intuitively, increasing dist leads to an over-inclusion of words as keywords, which results in a bias towards classifying predictions as trustworthy. This increases sensitivity while reducing specificity. Conversely, decreasing dist causes related keywords to be overlooked, leading to a bias towards classifications of predictions as untrustworthy, thereby reducing sensitivity and increasing specificity. Hence, we also address RQ2.1 by measuring the effectiveness of TOKI-lime with various values of dist using G-mean, a metric that balances sensitivity and specificity. Figure 9 reveals that the optimal value of dist ranges from 0.2 to 0.4, depending on the models under test. Across all trustworthiness datasets, when dist exceeds 0.6, G-mean drops significantly. In this case, TOKI considers all words as keywords and biases for labeling predictions as trustworthy. In the subsequent experiments, we set dist to 0.3, as it balances the effectiveness across the benchmark.
@@
RQ2.2: Comparing TOKI and Naive. TOKI consistently outperforms Naive across all datasets. Naive occasionally shows good precision, sensitivity and F1-score but relatively low specificity and G-mean, especially on Dong’s, Issues, and our trustworthiness datasets. This indicates that Naive is biased toward labeling predictions as trustworthy, resulting in misclassifying untrustworthy predictions as trustworthy.
@@
RQ2.4: The impact of explanation methods. We investigate this using Dong’s, CAMS, and HateXplain datasets, as other datasets only use LIME. Overall, TOKI demonstrates the best effectiveness with the SOTA LIME. TOKI-lime achieves higher accuracy and G-mean compared to TOKI-omis and TOKI-grad, while other metrics show no significant differences, except for the CAMS dataset. Explanation methods also impact TOKI’s efficiency. TOKI-lime is significantly slower due to LIME’s sampling of 5,000 neighbors, while omission and gradient are more lightweight. The model complexity also affects TOKI’s efficiency, as the fine-tuned models consume more time than the two simpler MLPs in Dong’s dataset.
@@
Table 4 compares the effectiveness of TOKI-guided attack method and A2T (Yoo and Qi, 2021). Overall, TOKI-guided attack method outperforms A2T, achieving a higher ASR and lower NP across all models. Adversarial examples generated by TOKI also have higher Bert and USE scores than those by A2T, except for the USE score on the ag_news dataset, demonstrating TOKI’s ability to preserve the semantic meaning of the original texts. Table 5 displays several adversarial examples generated by both methods. It is observed that TOKI can effectively find synonyms to attack with fewer perturbed words, while A2T requires more perturbations and sometimes fails. Interestingly, in Dong’s dataset, A2T performs nearly as well as the proposed method, likely because the models are simple MLPs and more easily attacked than transformer-based models in other datasets. However, in other datasets, the proposed method outperforms A2T by 2.6%–17.5% in ASR and requires 0.1–1.34 fewer NP. These results underscore the effectiveness of TOKI-guided attack method, particularly against models using SOTA architectures.
[Answer to RQ3] The adversarial attack method guided by TOKI outperforms the SOTA A2T. Our method achieves a higher success rate with fewer perturbations than A2T while preserving the semantic meaning of the original texts.
@@ 7.1. Implications
Prediction uncertainty and trustworthiness.
Experimental results show that highly confident predictions are more likely to be trustworthy. However, high confidence does not guarantee trustworthiness, and low confidence predictions can still be trustworthy. Relying on prediction uncertainty is ineffective for trustworthiness assessment, overlooking both trustworthy low and untrustworthy high confidence predictions. Prior work has a complex debate over the impact of prediction uncertainty, mainly on human trust in ML models. Several studies (Nguyen et al., 2015; Bussone et al., 2015) suggest that prediction uncertainty has a limited impact on human trust (Rechkemmer and Yin, 2022; Ovadia et al., 2019). Other studies (Zhang et al., 2020a) show that prediction certainty can improve human trust in ML models and increase the willingness to rely on high confidence predictions. However, we argue that whether prediction certainty increases human trust or not, it does not translate into improving the model’s trustworthiness. Model certainty can lead to overreliance on ML decisions without reflecting the true trustworthiness of the model. The model may make decisions based on spurious correlations, being confident in them rather than valid correlations in the training data. This fits the distinction between trust and trustworthiness defined by Kästner et al. (2021), where people can still trust an untrustworthy model.
Automated trustworthiness oracles.
TOKI identifies keywords for each class and uses them to measure the semantic relatedness between words and the class, effectively capturing the characteristics of classes. This makes TOKI outperform the baselines, including its variant that omits keywords to measure relatedness, which is biased for untrustworthy predictions.
@@
Implications for software engineering.
Trustworthiness is a crucial non-functional requirement of ML systems (Riccio et al., 2020). Trustworthiness oracles enable automating trustworthiness testing, which is particularly valuable in the iterative development of ML systems. They support continuous integration and delivery pipelines, monitoring in real-world SE environments, such as online testing or DevOps, and handling corner-cased and real-world inputs. Since TOKI is efficient and lightweight, it can be directly applied in online settings and real-world environments without extensive training, requiring only a one-time keyword identification. This makes TOKI a practical safety net for software systems that rely on ML text classification. TOKI’s outputs also provide actionable insights for feedback-driven improvement, reducing the need for human intervention.
Text classification is also a primary application in SE domains to automate tasks such as sentiment analysis of SE artifacts, including git commit comments, JIRA issues, and app reviews, assessing developers’ psychological states, analysing Q&A sites like StackOverflow, software requirements classification, and project issue categorisation. Integrating such oracles can foster greater trust in ML4SE systems and significantly advance their development and deployment.
@@ 7.2. Threats to Validity
To reduce threats to external validity, we evaluate TOKI on approximately 8,000 predictions across 11 models and datasets in various domains, including topic classification, sentiment analysis, clinical mental text classification, hate speech detection, and software issue classification. We also employ TOKI with various explanation methods to assess its performance.
Threats to construct validity can arise from the trustworthiness benchmark. We leverage existing studies, which do not directly label predictions as trustworthy or not, by measuring the explanations’ plausibility. To mitigate the threats, we collect an additional, more straightforward trustworthiness dataset. The bias in human annotations is addressed by combining all human-annotated trustworthiness labels for each prediction using plurality voting (Leon et al., 2017).
@@ 8.1. Machine Learning Testing
8.2. Explanation Method
Local explanations, in contrast to global counterparts, focus on how a specific input leads to a prediction. This makes them particularly suitable for understanding ML predictions. In text classification, various methods have been developed to generate local explanations.
Feature attribution-based explanations assess the relevance of individual input features, such as words, to a model’s prediction. Some (Li et al., 2016; Dong, 2018) achieve this by removing, masking, or modifying input features, and observing the impact on the prediction. Others calculate the gradient of the output with respect to the input (Mohebbi et al., 2021). Methods like LIME (Ribeiro et al., 2016) approximate the model’s behaviour with simpler, interpretable models.
Attention-based explanations leverage the attention mechanism, which often serves as a means to attend to the most relevant part of inputs (Wiegreffe and Pinter, 2019). Intuitively, they can capture meaningful correlations between intermediate states of the inputs, potentially explaining the model’s predictions (Zhao et al., 2024). Hence, many approaches (Yeh et al., 2024; Barkan et al., 2021) aim to explain the predictions solely based on the attention weights or by analysing the knowledge encoded in the attention.
Counterfactual explanations reveal model behavior by demonstrating how slight input changes impact outputs, highlighting key features influencing predictions (Zhao et al., 2024). Originally applied to ML tasks with explicit features and tabular datasets (Guidotti, 2024), these explanations have since been extended to other tasks, including text classification (Treviso et al., 2023).
While other local explanation methods are subject to extensive debate (Jain and Wallace, 2019), feature attribution explanations, such as the SOTA LIME (Ribeiro et al., 2016), have proven to be effective, faithful (Mariotti et al., 2024), and widely used by AI practitioners. Therefore, this paper focuses on feature attribution-based explanations. Exploring other local explanations for trustworthiness testing is a promising direction, which we leave for future work.
@@ 8.3. Making Use of Explanations
Explanations have served as a valuable tool for testing and improving ML models (Zhao et al., 2024). Various studies have used explanations to understand and debug models. Ribeiro et al. (2016) demonstrated the usefulness of local explanations in various tasks, such as model comparison and trust assessment. Lapuschkin et al. (2019) introduced a semi-automated approach that characterises and validates classification strategies. Thomas et al. (2019) used explanations to uncover input-prediction patterns. Several studies (Yoo and Qi, 2021; Zhang et al., 2020b) have applied explanations in adversarial attacks, primarily to determine word substitution order based on word importance. To the best of our knowledge, this paper introduces the first approach that leverages explanations to make perturbations, specifically, by weakening valid correlations and strengthening spurious ones. Explanations have also been integrated into the learning process to improve model performance and reliability. Ross et al. (2017) proposed an approach to discover multiple models for the same task with different classification strategies, allowing domain experts to choose the best one. Chen and Ji (2022) used explanations to improve the adversarial robustness of language models. Other studies (Schramowski et al., 2020; Ghai et al., 2021) leveraged explanations to make models right for the right scientific reasons. Similarly, Du et al. (2021) developed a framework to mitigate shortcuts, focusing on stop words, punctuation, and numbers. Recent studies (Linhardt et al., 2024) have also applied explanation regularisation to mitigate the impact of spurious correlations.
@@ 9. Conclusion
We investigate the trustworthiness oracle problem of text classifiers. Statistical evaluation reveals that while highly confident predictions are more likely to be trustworthy, some still lack trustworthiness due to reliance on spurious correlations. We propose TOKI, an automated trustworthiness oracle generation method. Experiments compare the effectiveness and efficiency of TOKI, TOKI’s ablation variant and the naive approach against the trustworthiness benchmark on ten models and datasets. Results show that TOKI outperforms other approaches. We also introduce a TOKI-guided adversarial attack method, which proves to be more effective than the SOTA A2T. In addition, several directions remain open for future work. We plan to explore alternative explanations, such as attention-based and counterfactual explanations, as well as other word embedding methods like sentence-transformers (Reimers and Gurevych, 2019). Further experiments on SE-specific datasets are also necessary to gain more insights into the trustworthiness problem in SE. Moreover, adapting TOKI to other data types, such as images and speech, presents a promising avenue in the future.
Data Availability.
Our data and replicate package are available at (Lam et al., 2024).